Pre-Algebra: The length of a rectangle is twice its width. The perimeter is 30 meters. Find the length and width.

a. length is 10 meters and the width is 5 meters
b. length is 7 meters and the width is 5 meters
c. length is 8 meters and the width is 7 meters
d. length is -10 meters and the width is -5 meters

Answer by Sumit56(2)

breadth or width = b
length= l = 2*b
Perimeter = 30 m
2 (l+b) = 30
l+b = 15
2b +b =15 (l=2b)
3b=15
b=5m
l= 2 * b = 2 * 5 = 10m
Right answer is a. l=10m and b=5m